OpenStudy (abbycross167):

What would the slope of the graph that goes through the two points (0, 4) and (-8, -4) be (show your work)? What would this graph look like? What should the equation of this line be? Can someone please help me figure out how to do this?

OpenStudy (greatlife44):

are you familiar with this formula? \[m = \frac{ y_{2}-y_{1} }{ x_{2}-x_{1} }\]

OpenStudy (abbycross167):

Yes ma'am/sir

OpenStudy (greatlife44):

any line has the form \[y = m(x)+b \] where m is the slope. the equation above helps you to find the slope which is the change in y with respect to x

OpenStudy (greatlife44):

so for our two points (0,4) (-8,-4) we must chose our x2 and y2

OpenStudy (greatlife44):

let's say x2 = -8 and y2 = -4 then x1 = 0 and y1 = 4 then we put this together \[\frac{ y_{2}-y_{1} }{ x_{2}-x_{1} } = \frac{ -4-4 }{ -8-0 } = \frac{ -8 }{ -8 } = 1\]

OpenStudy (greatlife44):

so our slope for this would be 1

OpenStudy (greatlife44):

we can go a step further with this too. we can actually construct this equation \[y = m(x)+b \] b = the y intercept, the point the graph touches the y axis when x = 0. they gave us (0,4). so our b = 4 and m =1 so the full equation is this. \[y = x+4 \]

OpenStudy (abbycross167):

thank you soo much!!! I'm sorry I didn't reply sooner my WiFi is slow